OpenStudy (anonymous):

simplify 9 (1/2) x 9 (1/2) using radical form PLEASE

OpenStudy (anonymous):

is this \[9^{\frac{1}{2}}\times 9^{\frac{1}{2}}\]

OpenStudy (anonymous):

is that 9^(1/2) * 9^(1/2)

OpenStudy (anonymous):

if so the answer is 9

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

but if you need to show your work in radical form you need to know: 9^(1/2) = \[\sqrt{9}\]